OpenStudy (anonymous):

square root 2 = 2cosx

OpenStudy (anonymous):

45 degree

OpenStudy (anonymous):

\[\sqrt{2}\]/2 = cosx x=cos\[\cos^{-1} (\sqrt{2}/2)\] x=45

OpenStudy (nice):

cosx = square root 2 / 2 x = arc cos(square root 2 / 2)

OpenStudy (anonymous):

Thank you